(x+5)^2/3=4
i don't have a clue.

Answer by checkley77(12844)

(x+5)^2/3=4
the exponent (^2/3) is another way of representin the cuberoot of the square of the term (x+5).
cubert(x+5)^2=4
now if we cube each side of the equation we get:
(x+5)^2=4^3
x^2+10x+25=64
x^2+10x+25-64=0
x^2+10x-39=0
(x+13)(x-3)=0
x+13=0
x=-13 answer.
x-3=0
x=3 answer.
proof:
cubert(3+5)^2=4
cubert(8^2)=4
cubert64=4
4=4